OpenStudy (waheguru):

plz help need help fast -(6-y)(-2)

OpenStudy (anonymous):

12-2y

OpenStudy (waheguru):

plzz stay 1 more questionm

OpenStudy (waheguru):

once u multiply the -1 it becaimes -6+y*-2 and then -6-2y ??????

OpenStudy (anonymous):

-(6-y)*(-2) = (-1)*(6-y)*(-2) = (-6+y)*(-2) = 12- 2y

OpenStudy (anonymous):

you keep the brackets and multiply everything by -2

OpenStudy (anonymous):

2(6-y)